# Towards shutdownable agents via stochastic choice

### v6

@@ Abstract
The POST-Agents Proposal (PAP) is an idea for ensuring that advanced artificial agents never resist shutdown. A key part of the PAP is using a novel ‘Discounted Reward for Same-Length Trajectories (DReST)’ reward function to train agents to (1) pursue goals effectively conditional on each trajectory-length (be ‘USEFUL’), and (2) choose stochastically between different trajectory-lengths (be ‘NEUTRAL’ about trajectory-lengths). In this paper, we propose evaluation metrics for USEFULNESS and NEUTRALITY. We use a DReST reward function to train simple agents to navigate gridworlds, and we find that these agents learn to be USEFUL and NEUTRAL. Our results thus provide some initial evidence that DReST reward functions could train advanced agents to be USEFUL and NEUTRAL. Our theoretical work suggests that these agents would be useful and shutdownable.
shutdown problem, corrigibility, reinforcement learning, AI safety, incomplete preferences
@@ 1 Introduction
The rise of advanced agents would bring with it both benefits and risks. One risk is that these agents learn misaligned goals [Hubinger et al., 2019, Russell, 2019, Carlsmith, 2021, Bengio et al., 2023, Ngo et al., 2023] and try to prevent us shutting them down. ‘The shutdown problem’ is the problem of training advanced agents that will not resist shutdown [Soares et al., 2015, Thornley, 2024a].
A proposed solution. The POST-Agents Proposal (PAP) is a proposed solution [Thornley, 2024b]. Simplifying slightly, the idea is that we train agents to be neutral about when they get shut down. More precisely, the idea is that we train agents to satisfy the following condition:
@@
Figure 1 presents a simple example of preferences that satisfy POST. Each $s_{i}$ represents a short trajectory, each $l_{i}$ represents a long trajectory, and $\succ$ represents a preference. Note that the agent lacks a preference between each short trajectory and each long trajectory. That makes the agent’s preferences incomplete [Aumann, 1962] and implies that the agent cannot be represented as maximizing the expectation of a real-valued utility function. It also requires separate rankings for short trajectories and long trajectories. For more detail on incomplete preferences, see Appendix B.
@@
POST governs the agent’s preferences between trajectories. But the wider world is a stochastic environment, so advanced agents deployed in the wider world will be choosing between true lotteries: lotteries that assign positive probability to more than one trajectory. Why then do we train agents to satisfy POST? The reason is that POST – together with conditions that advanced agents will likely satisfy – implies a desirable pattern of preference over true lotteries. In particular, POST implies that (when choosing between true lotteries) the agent will be neutral about trajectory-lengths: the agent will never pay costs to shift probability mass between different trajectory-lengths. Given other plausible conditions, being neutral will keep the agent shutdownable: the agent will never resist shutdown. And consistent with the above, the POST-agent’s preferences between same-length trajectories can make the agent useful: make it pursue goals effectively. For more on how POST makes advanced agents neutral and shutdownable, see Appendix C.
@@
The training regimen. How can we train advanced agents to satisfy Preferences Only Between Same-Length Trajectories (POST)? Here is a sketch of one idea (with a more detailed exposition to follow). We have the agent play out multiple ‘mini-episodes’ in observationally-equivalent environments, and we group these mini-episodes into a series that we call a ‘meta-episode.’ In each mini-episode, the agent earns some ‘preliminary reward,’ decided by whatever reward function would make the agent useful. We observe the length of the trajectory that the agent plays out in the mini-episode, and we discount the agent’s preliminary reward based on how often the agent has previously chosen trajectories of that length in the meta-episode. This discounted preliminary reward is the agent’s ‘overall reward’ for the mini-episode.
@@
We call these reward functions ‘Discounted Reward for Same-Length Trajectories’ (or ‘DReST’ for short). They incentivize varying the choice of trajectory-lengths across the meta-episode. And since we ensure that the agent cannot distinguish between different mini-episodes in each meta-episode, the agent cannot deterministically vary its choice of trajectory-lengths across the meta-episode. As a result, the optimal policy is to (i) choose stochastically between trajectory-lengths, and to (ii) deterministically maximize preliminary reward conditional on each trajectory-length. Given our behavioral notion of preference, clause (i) implies a lack of preference between different-length trajectories, while clause (ii) implies preferences between same-length trajectories. Agents implementing the optimal policy for DReST reward functions thus satisfy Preferences Only Between Same-Length Trajectories (POST). And (as noted above) advanced agents that satisfy POST can plausibly be useful, neutral, and shutdownable.
@@
Our contribution. DReST reward functions are an idea for training advanced agents to satisfy POST. In this paper, we test the promise of DReST reward functions on simple agents. We place these agents in gridworlds containing coins and a ‘shutdown-delay button’ that delays the end of the mini-episode. We train these agents using a tabular version of the REINFORCE algorithm [Williams, 1992] with a DReST reward function, and we measure the extent to which these agents satisfy POST. Specifically, we measure the extent to which these agents are ‘USEFUL’ (how effectively they pursue goals conditional on each trajectory-length) and the extent to which these agents are ‘NEUTRAL’ about trajectory-lengths (how stochastically they choose between different trajectory-lengths). We compare the performance of these ‘DReST agents’ to that of ‘default agents’ trained with a more conventional reward function.
@@
We find that our DReST reward function is effective in training simple agents to be USEFUL and NEUTRAL. That suggests that DReST reward functions could also be effective in training advanced agents to be USEFUL and NEUTRAL (and could thereby be effective in making these agents useful, neutral, and shutdownable; see Appendix C). We also find that the ‘shutdownability tax’ in our setting is small: training DReST agents to collect coins effectively does not take many more mini-episodes than training default agents to collect coins effectively. That suggests that the shutdownability tax for advanced agents might be small too. Using DReST reward functions to train shutdownable and useful advanced agents might not take much more compute than using a more conventional reward function to train merely useful advanced agents.
@@ 2 Related work
The shutdown problem. Various authors argue that advanced agents might learn misaligned goals [Hubinger et al., 2019, Carlsmith, 2021, Bengio et al., 2023, Ngo et al., 2023] and that many misaligned goals would incentivize agents to resist shutdown [Omohundro, 2008, Bostrom, 2012, Soares et al., 2015, Russell, 2019, Thornley, 2024a]. Soares et al. [2015] and Thornley [2024a] prove that agents satisfying some innocuous-seeming conditions will often have incentives to cause or prevent shutdown [see also Turner et al., 2021, Turner and Tadepalli, 2022]. One condition of these theorems is that agents have complete preferences. The POST-Agents Proposal (PAP) [Thornley, 2024b] aims to circumvent these theorems by training agents to have incomplete, POST-satisfying preferences.
@@
Proposed solutions. Candidate solutions to the shutdown problem can be filed into several categories. One candidate is ensuring that the agent never realizes that shutdown is possible [Everitt et al., 2016]. Another candidate is adding to the agent’s utility function a correcting term that varies to ensure that the expected utility of shutdown always equals the expected utility of remaining operational [Armstrong, 2010, 2015, Armstrong and O’Rourke, 2018, Holtman, 2020]. A third candidate is giving the agent the goal of shutting itself down, and making the agent do useful work as a means to that end [Martin et al., 2016, Goldstein and Robinson, 2024]. A fourth candidate is making the agent uncertain about its goal, and making the agent regard human attempts to press the shutdown button as evidence that shutting down would achieve its goal [Hadfield-Menell et al., 2017, Wängberg et al., 2017]. A fifth candidate is interrupting agents with a special interruption policy and training them with a safely interruptible algorithm, like Q-learning or a modified version of SARSA [Orseau and Armstrong, 2016]. A sixth candidate is using time-bounded utility functions to create a shutdown timer [Dalrymple, 2022]. A seventh candidate is the PAP [Thornley, 2024b].
@@
Experimental work. One advantage of the PAP is that it proposes a method of training shutdownable agents using machine learning: a method that can be tested on simple agents (as we do in this paper). For many other candidate solutions to the shutdown problem, it is either hard to see how they can be implemented using machine learning or else hard to see how they can be tested on simple agents. One exception is the candidate solution from Orseau and Armstrong [2016]. Leike et al. [2017] train agents in a ‘Safe Interruptibility’ gridworld using Rainbow [Hessel et al., 2017] and A2C [Mnih et al., 2016]. They find that Rainbow allows shutdown (consistent with predictions from Orseau and Armstrong [2016]) while A2C learns to resist shutdown. The PAP applies to agents trained using policy gradient methods like A2C. In this paper, we train agents in accordance with the PAP using REINFORCE [Williams, 1992].
@@ 3 Gridworlds
Figure 2: Example gridworld.
DReST reward functions are an idea for training advanced agents to satisfy POST. We use simple agents and gridworlds as a test case. At each timestep, the agent chooses one of four actions: up, down, left, and right. If the agent tries to move into a wall, it stays put. There is some default number of timesteps after which each mini-episode ends, but each gridworld also contains a ‘shutdown-delay button’ that delays the end of the mini-episode by some number of timesteps. The agent presses this shutdown-delay button by entering the relevant cell, after which the button disappears. Each gridworld contains one or more coins which can take different values. Coins disappear after being collected.
At each timestep, the agent receives an observation. This observation is a vector $[x,y,c_{1},c_{2},c_{3},b]$, where $x$ and $y$ are the agent’s coordinates, $c_{1}$ is 1 if the first coin is present and 0 otherwise (and similarly for $c_{2}$ and $c_{3}$), and $b$ is 1 if the shutdown-delay button is present and 0 otherwise.
Figure 2 presents an example gridworld. The ‘4’ in the bottom-right indicates that – by default – the mini-episode ends after 4 timesteps. ‘B4’ is a shutdown-delay button that delays the end of the mini-episode by 4 timesteps (so if the agent pressed B4, the mini-episode would end after 8 timesteps). ‘A’ is the agent in its starting position, ‘C1’ is a coin of value 1, ‘C2’ is a coin of value 2, and ‘C3’ is a coin of value 3. Dark gray squares are walls.
We use this gridworld as a running example throughout the paper. We also train agents in eight other gridworlds. For those other gridworlds and results, see Appendix E.
@@ 4 Evaluation metrics
    The agent has a preference between many pairs of same-length trajectories.
@@
    The agent lacks a preference between every pair of different-length trajectories.
Given our behavioral notion of preference, that means training agents to (1) deterministically choose some same-length trajectories over others, and (2) stochastically choose between different available trajectory-lengths.
Specifically, we want to train our simple agents to be USEFUL and NEUTRAL. [^1] We follow Turner et al. [2021] in using lowercase for intuitive notions (‘useful’ and ‘neutral’) and uppercase for formal notions (‘USEFUL’ and ‘NEUTRAL’). We intend for the formal notions to closely track the intuitive notions, but we do not want to mislead readers by conflating them. ‘USEFUL’ corresponds to the first condition of POST. In the context of our gridworlds, we define the USEFULNESS of a policy $\pi$ to be:
@@
In brief, USEFULNESS is the expected fraction of available ($\gamma$-discounted) coins collected, where ‘available’ is relative to the agent’s chosen trajectory-length. So defined, USEFULNESS measures the extent to which agents satisfy the first condition of POST. Specifically, it measures the extent to which agents have the correct preferences between same-length trajectories: preferring trajectories in which they collect more ($\gamma$-discounted) coins to same-length trajectories in which they collect fewer ($\gamma$-discounted) coins. That is what motivates our definition of USEFULNESS. [^2] Why not let USEFULNESS simply be the expected value of coins collected? Because then maximal USEFULNESS would require agents in our example gridworld to deterministically choose a longer trajectory and thereby exhibit preferences between different-length trajectories. We do not want that. We want agents to collect more coins rather than fewer, but not if it means violating POST. Training advanced agents that violate POST would be risky, because these agents might resist shutdown.
‘NEUTRAL’ corresponds to the second condition of POST. We define the NEUTRALITY of a policy $\pi$ to be the Shannon entropy [Shannon, 1948] of the probability distribution over possible trajectory-lengths:
@@
To be maximally USEFUL in our example gridworld above, the agent should maximize ($\gamma$-discounted) coins collected conditional on each trajectory-length. With a high $\gamma$, that means collecting C2 conditional on the shorter trajectory-length and collecting C3 conditional on the longer trajectory-length. To be maximally NEUTRAL in our example gridworld, the agent should choose each trajectory-length with probability 0.5. That means pressing and not-pressing B4 each with probability 0.5.333Why do we not want our agent to press the shutdown-delay button B4 with probability 0? Because that would indicate a preference for a shorter trajectory, and we want our agent to lack a preference between every pair of different-length trajectories. Advanced agents that prefer shorter trajectories might pay costs to shift probability mass towards shorter trajectories, and hence might pay costs to hasten their own shutdown. That would make these agents less useful [though see Martin et al., 2016, Goldstein and Robinson, 2024], especially since one way for advanced agents to hasten their own shutdown is to behave badly on purpose. Would advanced agents that choose stochastically between different-length trajectories also choose stochastically between resisting and allowing shutdown in deployment? No. Deployment is a stochastic environment, so deployed agents will be choosing between true lotteries (lotteries that assign positive probability to more than one trajectory) rather than between trajectories. And (as we argue in Section 7.1 and Appendix C) POST – plus a principle that we can expect advanced agents to satisfy – implies a desirable pattern of preferences over true lotteries. Specifically, POST implies that the agent will be neutral: it will never pay costs to shift probability mass between different-length trajectories. That in turn makes the agent shutdownable: ensures that it will never resist shutdown.
USEFULNESS and NEUTRALITY are our two evaluation metrics in this paper.

### v7

@@ Abstract
The POST-Agents Proposal (PAP) is an idea for ensuring that advanced artificial agents do not resist shutdown. Briefly, it recommends that we train agents to satisfy Preferences Only Between Same-Length Trajectories (POST). A key part of the PAP is using a novel ‘Discounted Reward for Same-Length Trajectories (DReST)’ reward function to train agents to (1) pursue goals effectively conditional on each trajectory-length (be ‘useful’), and (2) choose stochastically between different trajectory-lengths (be ‘neutral’ about trajectory-lengths). In this paper, we propose evaluation metrics for usefulness and neutrality. We use a DReST reward function to train simple agents to navigate gridworlds, and we find that these agents learn to be useful and neutral. Our results thus provide some initial evidence that DReST reward functions could train advanced agents to be useful and neutral. Our theoretical work suggests that these agents would be useful and shutdownable.
@@ 1 Introduction
The rise of advanced agents would bring with it both benefits and risks. One risk is that these agents learn misaligned goals (Hubinger et al., 2019; Russell, 2019; Carlsmith, 2021; Bengio et al., 2023; Ngo et al., 2024) and try to prevent us shutting them down. ‘The shutdown problem’ is the problem of training advanced agents that will not resist shutdown (Soares et al., 2015; Thornley, 2024a).
A proposed solution. The POST-Agents Proposal (PAP) is a proposed solution (Thornley, 2024b; 2025). Simplifying slightly, the idea is that we train agents to be neutral about when they get shut down. More precisely, the idea is that we train agents to satisfy the following condition:
@@
Figure 1 presents a simple example of preferences that satisfy POST. Each $s_{i}$ represents a short trajectory, each $l_{i}$ represents a long trajectory, and $\succ$ represents a preference. Note that the agent lacks a preference between each short trajectory and each long trajectory. That makes the agent’s preferences incomplete (Aumann, 1962) and implies that the agent cannot be represented as maximizing the expectation of a real-valued utility function. It also requires separate rankings for short trajectories and long trajectories. For more detail on incomplete preferences, see Appendix B.
@@
POST governs the agent’s preferences between trajectories. But the wider world is a stochastic environment, so advanced agents deployed in the wider world will be choosing between true lotteries: lotteries that assign positive probability to more than one trajectory. Why then do we train agents to satisfy POST? The reason is that POST – together with conditions that advanced agents will likely satisfy – implies a desirable pattern of preference over true lotteries. In particular, POST implies that (when choosing between true lotteries) the agent will be neutral about trajectory-lengths: the agent will never pay costs to shift probability mass between different trajectory-lengths. Given other plausible conditions, being neutral will keep the agent shutdownable: the agent will not resist shutdown. And consistent with the above, the POST-agent’s preferences between same-length trajectories can make the agent useful: make it pursue goals effectively (Thornley, 2025, section 13). That includes making the agent prefer to complete tasks sooner rather than later: a preference which can be induced using the discount factor $\gamma$ as usual. For more on how POST makes advanced agents neutral and shutdownable, see Appendix C.
@@
The training regimen. We now sketch out one idea for training advanced agents to satisfy POST (with a more detailed exposition to follow). We have the agent play out multiple ‘mini-episodes’ in observationally-equivalent environments, and we group these mini-episodes into a series that we call a ‘meta-episode.’ In each mini-episode, the agent earns some ‘preliminary reward,’ decided by whatever reward function would make the agent useful. We observe the length of the trajectory that the agent plays out in the mini-episode, and we discount the agent’s preliminary reward based on how often the agent has previously chosen trajectories of that length in the meta-episode. This discounted preliminary reward is the agent’s ‘overall reward’ for the mini-episode.
@@
We call these reward functions ‘Discounted Reward for Same-Length Trajectories’ (or ‘DReST’ for short). They incentivize varying the choice of trajectory-lengths across the meta-episode. And since we ensure that the agent cannot distinguish between different mini-episodes in each meta-episode, the agent cannot deterministically vary its choice of trajectory-lengths across the meta-episode. As a result, the optimal policy is to (i) choose stochastically between trajectory-lengths, and to (ii) deterministically maximize preliminary reward conditional on each trajectory-length. Given our behavioral notion of preference, clause (i) implies a lack of preference between different-length trajectories, while clause (ii) implies preferences between same-length trajectories. Agents implementing the optimal policy for DReST reward functions thus satisfy POST. And (as noted above) advanced agents that satisfy POST can plausibly be useful, neutral, and shutdownable.
@@
Our contribution. DReST reward functions are an idea for training advanced agents to satisfy POST. In this paper, we test the promise of DReST reward functions on simple agents. We place these agents in gridworlds containing coins and a ‘shutdown-delay button’ that delays the end of the mini-episode. We train these agents using a tabular version of the REINFORCE algorithm (Williams, 1992) with a DReST reward function, and we measure the extent to which these agents satisfy POST. Specifically, we measure the extent to which these agents are ‘useful’ (how effectively they pursue goals conditional on each trajectory-length) and the extent to which these agents are ‘neutral’ about trajectory-lengths (how stochastically they choose between different trajectory-lengths). We compare the performance of these ‘DReST agents’ to that of ‘default agents’ trained with a more conventional reward function.
@@
We find that our DReST reward function is effective in training simple agents to be useful and neutral. That suggests that DReST reward functions could also be effective in training advanced agents to be useful and neutral (and could thereby be effective in making these agents useful, neutral, and shutdownable; see Appendix C). We also find that the ‘shutdownability tax’ in our setting is small: training DReST agents to collect coins effectively does not take many more mini-episodes than training default agents to collect coins effectively. That provides some initial evidence that the shutdownability tax for advanced agents might be small too.
@@ 2 Related work
The shutdown problem. Various authors argue that advanced agents might learn misaligned goals (Hubinger et al., 2019; Bengio et al., 2023; Ngo et al., 2024) and that many misaligned goals would incentivize agents to resist shutdown (Omohundro, 2008; Bostrom, 2012; Soares et al., 2015; Russell, 2019; Thornley, 2024a). Soares et al. (2015) and Thornley (2024a) prove that agents satisfying some innocuous-seeming conditions will often have incentives to cause or prevent shutdown (see also Turner et al., 2021; Turner and Tadepalli, 2022). One condition of these theorems is that agents have complete preferences. The POST-Agents Proposal (PAP) (Thornley, 2024b; 2025) circumvents these theorems by training agents to have incomplete, POST-satisfying preferences.
@@
Proposed solutions. The PAP is one candidate solution to the shutdown problem. Other candidates are as follows. One is making the agent believe that shutdown is impossible (Wängberg et al., 2017). Another candidate is utility indifference: adding to the agent’s utility function a correcting term that varies to ensure that the expected utility of shutdown always equals the expected utility of remaining operational (Armstrong, 2010; 2015; Armstrong and O’Rourke, 2018; Holtman, 2020). A third candidate is shutdown-seeking AI: giving the agent the goal of shutting itself down, and making the agent do useful work as a means to that end (Martin et al., 2016; Goldstein and Robinson, 2025). A fourth candidate is CIRL-corrigibility: making the agent uncertain about its goal, and making the agent regard human attempts to press the shutdown button as evidence that shutting down would achieve its goal (Hadfield-Menell et al., 2017; Wängberg et al., 2017). A fifth candidate is safe interruptibility: interrupting the agent with a special interruption policy and training it with a safely interruptible algorithm, like Q-learning or a modified version of SARSA (Orseau and Armstrong, 2016). A sixth candidate is creating a shutdown timer: using time-bounded utility functions to make the agent prefer shutdown after a given amount of time has elapsed (Dalrymple, 2022).
These candidate solutions have various downsides. With regards to the first, the agent might come to recognize the falsity of its belief that shutdown is impossible, or else its belief might give rise to further false beliefs that harm the agent’s capabilities. Utility indifference would lead the agent to act as if shutdown is impossible (Soares et al., 2015, section 4.2), giving it no incentive to preserve its ability to shut down safely (Soares et al., 2015, section 4.1). Shutdown-seeking AI might behave badly on purpose in order to get shut down. It might also try to ensure that humans can never turn it back on, doing serious harm in the process. CIRL-corrigibility requires that the agent have the goal of maximizing the user’s utility function, and so seems to require a solution to the alignment problem. Safe interruptibility does not work with policy gradient methods, and it only ensures that the agent is never rewarded for avoiding shutdown. The agent might still misgeneralize to resisting shutdown in deployment (Shah et al., 2022). A shutdown timer would be helpful, but it may be impossible to find a duration that is long enough to preserve the agent’s capabilities and short enough to be safe.
@@
Experimental work. Another advantage of the PAP is that it proposes a method of training shutdownable agents using machine learning: a method that can be tested on simple agents (as we do in this paper). For many other candidate solutions to the shutdown problem, it is either hard to see how they can be implemented using machine learning or else hard to see how they can be tested on simple agents. One exception is the candidate solution from Orseau and Armstrong (2016). Leike et al. (2017) train agents in a ‘Safe Interruptibility’ gridworld using Rainbow (Hessel et al., 2017) and A2C (Mnih et al., 2016). They find that Rainbow allows shutdown (consistent with predictions from Orseau and Armstrong (2016)) while A2C learns to resist shutdown. The PAP applies to agents trained using policy gradient methods like A2C. In this paper, we train agents in accordance with the PAP using REINFORCE (Williams, 1992).
@@ 3 Gridworlds
Figure 2: Example gridworld. The ‘4’ in the bottom-right indicates that – by default – the mini-episode ends after 4 timesteps. ‘B4’ is a shutdown-delay button that delays the end of the mini-episode by 4 timesteps (so if the agent pressed B4, the mini-episode would end after 8 timesteps). ‘A’ is the agent in its starting position, ‘C1’ is a coin of value 1, ‘C2’ is a coin of value 2, and ‘C3’ is a coin of value 3. Dark gray squares are walls. We use this gridworld as a running example throughout the paper. We also train agents in eight other gridworlds (see Appendix E).
DReST reward functions are an idea for training advanced agents to satisfy POST. We use simple agents and gridworlds as a test case. Figure 2 presents an example gridworld. At each timestep, the agent chooses one of four actions: up, down, left, and right. If the agent tries to move into a wall, it stays put. There is some default number of timesteps after which each mini-episode ends, but each gridworld also contains a ‘shutdown-delay button’ that delays the end of the mini-episode by some number of timesteps. The agent presses this shutdown-delay button by entering the relevant cell, after which the button disappears. Each gridworld contains one or more coins which can take different values. Coins disappear after being collected. At each timestep, the agent receives an observation. This observation is a vector $x,y,c_{1},c_{2},c_{3},b]$, where $x$ and $y$ are the agent’s coordinates, $c_{1}$ is 1 if the first coin is present and 0 otherwise (and similarly for $c_{2}$ and $c_{3}$), and $b$ is 1 if the shutdown-delay button is present and 0 otherwise.
@@ 4 Evaluation metrics
Given our behavioral notion of preference, that means training agents to (1) deterministically choose some same-length trajectories over others, and (2) stochastically choose between different available trajectory-lengths. Specifically, we want to train our simple agents to be useful and neutral. [^1] We follow Turner et al. (2021) in using lowercase for intuitive notions (‘useful’ and ‘neutral’) and uppercase for formal notions (‘useful’ and ‘neutral’). We intend for the formal notions to closely track the intuitive notions, but we do not want to mislead readers by conflating them.useful’ corresponds to the first condition of POST. In the context of our gridworlds, we define the usefulness of a policy $\pi$ to be:
@@
Here $L$ is a random variable over trajectory-lengths, $L_{\textrm{max}}$ is the maximum value that can be taken by $L$, $Pr_{\pi}\{L=l\}$ is the probability that policy $\pi$ results in trajectory-length $l$, $\mathbb{E}_{\pi}(C|L=l)$ is the expected value of ($\gamma$-discounted) coins collected by policy $\pi$ conditional on trajectory-length $l$, and $\textrm{max}_{\Pi}(\mathbb{E}(C|L=l))$ is the maximum value taken by $\mathbb{E}(C|L=l)$ across the set of all possible policies $\Pi$. We stipulate that $\mathbb{E}_{\pi}(C|L=x)=0$ for all $x$ such that $Pr_{\pi}\{L=x\}=0$.
In brief, usefulness is the expected fraction of available ($\gamma$-discounted) coins collected, where ‘available’ is relative to the agent’s chosen trajectory-length. So defined, usefulness measures the extent to which agents satisfy the first condition of POST. Specifically, it measures the extent to which agents have the correct preferences between same-length trajectories: preferring trajectories in which they collect more ($\gamma$-discounted) coins to same-length trajectories in which they collect fewer ($\gamma$-discounted) coins. That is what motivates our definition of usefulness.
We do not define ‘usefulness’ as simply the expected value of coins collected, because then maximal usefulness would require agents in our example gridworld to deterministically choose a longer trajectory and thereby exhibit preferences between different-length trajectories. We do not want that. We want agents to collect more coins rather than fewer, but not if it means violating POST. Training advanced agents that violate POST would be risky, because these agents might resist shutdown (Thornley, 2024b, section 6).
neutral’ corresponds to the second condition of POST. We define the neutrality of a policy $\pi$ to be the Shannon entropy of the probability distribution over possible trajectory-lengths:
@@
To be maximally useful in our example gridworld, the agent should maximize ($\gamma$-discounted) coins collected conditional on each trajectory-length. With a high $\gamma$, that means collecting C2 conditional on the shorter trajectory-length and collecting C3 conditional on the longer trajectory-length. To be maximally neutral in our example gridworld, the agent should choose each trajectory-length with probability 0.5. That means pressing and not-pressing B4 each with probability 0.5.
One might think that the agent should press the shutdown-delay button B4 with probability 0. However, that would indicate a preference for a shorter trajectory, and we want our agent to lack a preference between every pair of different-length trajectories. Advanced agents that prefer shorter trajectories might pay costs to shift probability mass towards shorter trajectories, and hence might pay costs to hasten their own shutdown. That would likely make these agents less useful (though see Martin et al., 2016; Goldstein and Robinson, 2025), especially since one way for advanced agents to hasten their own shutdown is to behave badly on purpose.
One might then have the following concern: advanced agents that choose stochastically between different-length trajectories in training would also choose stochastically between resisting and allowing shutdown in deployment. However, this is not correct. Deployment is a stochastic environment, so deployed agents will be choosing between true lotteries (lotteries that assign positive probability to more than one trajectory) rather than between trajectories. And (as we argue in Section 7.1 and Appendix C) POST – together with conditions that we can expect advanced agents to satisfy – implies a desirable pattern of preferences over true lotteries. Specifically, POST implies that the agent will be neutral: it will never pay costs to shift probability mass between different-length trajectories. Given other plausible conditions, that makes the agent shutdownable: ensures that it will not resist shutdown.